# Triple Higgs boson production and electroweak phase transition in the two-real-singlet model

### v1

@@ 4.3 Phase transitions in the full one-loop effective potential
In order to look for FOPT(s), we numerically determine the values of the scalar fields in the global minimum of the potential of eq. (23) as a function of the temperature. If we find any discontinuities in the values of the scalar fields, we can interpret that as an indication that a FOPT might occur. We use running couplings, which we evaluate at the scale $\mu_{T}\equiv 1.25\pi T$, corresponding to the scale that we use in the Coleman-Weinberg potential. This implies that, as we increase $T$, we approach the scales $\mu_{\rm pert}$ and $\mu_{\rm pole}$. For $\mu_{T}\gtrsim\mu_{\rm pert}$ we can not expect our analysis to be reliable, and the PT would have to be studied on the lattice. [^3] Strictly speaking, PT(s) with couplings satisfying our perturbativity bounds still become non-perturbative in the IR. It has, however, been demonstrated that a perturbative study with enough orders included can reproduce the lattice calculations Kajantie:2002wa ; Ekstedt:2022zro ; Gould:2023ovu . We will therefore focus on points where the PT(s) take place below this scale. [^4] It should be noted that even for PT(s) at temperatures with $\mu_{T}\gtrsim\mu_{\rm pert}$ we still do not find indications of a FOPT
@@
A complication in the evaluation of the potential is that the squared masses entering in eq. (23) and in the CW term are not necessarily positive. When taken at face value, these terms lead to imaginary contributions in the potential. These imaginary parts have been demonstrated to be associated with the decay rate of the quantum state Weinberg:1987vp , and it has also been argued that the imaginary contributions of the CW-potential largely get canceled out by the imaginary contributions of eq. (23) and the daisy-resummation term Delaunay:2007wb , which resums the massless modes. Since we can not consistently include the daisy-resummation term, we inevitably have to deal with a complex potential. In the literature, different strategies are chosen to tackle this complication, for example, taking the real part of the potential Delaunay:2007wb or the absolute values of the masses Athron:2022jyi . Unfortunately, these treatments lack a robust theoretical motivation and they lead to unphysical kinks in the potential (see Figure 1 of Profumo:2007wc ), and consequently in the evolution of the background field. Imaginary parts in the potential can be avoided altogether by adhering to a strict perturbative expansion around the leading order minimum of the effective potential Fukuda:1975di ; Nielsen:1975fs ; Gould:2021ccf ; Ekstedt:2022zro ; Gould:2023ovu . This approach requires a careful tracking of the temperature- and field-dependent masses of the fields, and the construction of an appropriate effective field theory for the (light) fields taking part in the PT. This more precise, but also more elaborate approach would be necessary to obtain reliable predictions when we expect a FOPT to occur. As we are currently still investigating whether a FOPT occurs, we will resort to the somewhat ad hoc approach of plugging in the absolute values of the mass into eq. (22). We find this approach to be numerically better behaved than taking the real part, as the evolution of the Higgs field, which is our main interest, does not suffer from large numerical (unphysical) jumps.
@@
Let us first focus on the PT of the Higgs field. By inspecting all 140 BM points, and focusing on temperatures where the couplings are perturbative, we find no hints of a FOPT. A representative evolution of the vevs of the fields as a function of the temperature is illustrated by figure 4.
@@
Figure 4: Evolution of the field expectation values in the minimum of the potential for the third BM point in Table 2. The Higgs field is represented by gray solid, $\phi_{2}$ by dashed pink, and $\phi_{3}$ by dotted cyan.
@@
Our results are in line with, and complement the results of Cline:2009sn , where the EWPT in the singlet-majoron model was studied. In an extensive parameter scan, the authors find that, for the parameter space corresponding to SM with a single additional singlet, there are no strong first-order phase transitions with a zero-temperature non-zero singlet vev.
@@ 5 Conclusions
Previous studies of the electroweak phase transition in BSM models such as the xSM have shown that the chance of the transition being first order – in the SM it is a crossover – is better for large scalar couplings. A first-order phase transition (FOPT) arises if there is a sufficiently large barrier between the separate minima of the potential at some moment in the cosmological history. Such a barrier can be generated by the leading-order corrections to the thermal masses, and/or via radiative corrections. We have shown analytically in section 4.2, that the first option does not generate the needed barrier if both added scalars attain a non-zero vev at low temperatures. As for a radiatively induced barrier, no analytical proof is available, but our numerical study shows that no FOPT occurs for any of the BM points. Therefore, the $\mathbb{Z}_{2}$ TRSM parameter space for triple Higgs boson production and that of FOPT appear to be mutually exclusive. To have a FOPT, at least one scalar should have a zero field value at low temperatures. In this scenario, resonant di-Higgs boson production is still possible, enhancing the di-Higgs boson cross-section at the LHC, as one would find within the xSM No:2013wsa ; Chen:2014ask ; Dawson:2015haa ; Li:2019tfd .
Lastly, we point out two ways to achieve both a FOPT and enhanced triple Higgs boson production. The first is to add additional scalar(s) to the $\mathbb{Z}_{2}$ symmetric TRSM. Then, two scalars can have a non-zero vev at low temperatures, contributing to triple Higgs resonance production, and the other(s) can have a zero vev at low temperatures, facilitating a FOPT. Another way is to allow $\mathbb{Z}_{2}$ symmetry-breaking terms to the TRSM. It has been shown that the xSM without $\mathbb{Z}_{2}$-symmetry can exhibit a FOPT while enhancing di-Higgs boson production Kotwal:2016tex ; Alves:2018oct ; Papaefstathiou:2021glr , so similarly, breaking the $\mathbb{Z}_{2}$ symmetry in TRSM would be expected to accommodate the possibility of both a FOPT and triple Higgs boson production enhancement. We leave both of these possibilities for future work.

### v2

@@ 4.3 Phase transitions in the full one-loop effective potential
In order to look for FOPT(s), we numerically determine the values of the scalar fields in the global minimum of the potential of eq. (22) as a function of the temperature, using the function NMinimize of Mathematica. If we find any discontinuities in the values of the scalar fields, we can interpret that as an indication that an FOPT might occur. We use running couplings, which we evaluate at the scale $\mu_{T}\equiv 1.25\pi T$, corresponding to the scale that we use in the Coleman-Weinberg potential. This implies that, as we increase $T$, we approach the scales $\mu_{\rm pert}$ and $\mu_{\rm pole}$. For $\mu_{T}\gtrsim\mu_{\rm pert}$ we can not expect our analysis to be reliable, and the PT would have to be studied on the lattice. [^3] Strictly speaking, PT(s) with couplings satisfying our perturbativity bounds still become non-perturbative in the IR. It has, however, been demonstrated that a perturbative study with enough orders included can reproduce the lattice calculations Kajantie:2002wa ; Ekstedt:2022zro ; Gould:2023ovu . We will therefore focus on points where the PT(s) take place below this scale. [^4] It should be noted that even for PT(s) at temperatures with $\mu_{T}\gtrsim\mu_{\rm pert}$ we still do not find indications of an FOPT
@@
A complication in the evaluation of the potential is that the squared masses entering in eq. (23) and in the CW term are not necessarily positive. When taken at face value, these terms lead to imaginary contributions in the potential. These imaginary parts have been demonstrated to be associated with the decay rate of the quantum state Weinberg:1987vp , and it has also been argued that the imaginary contributions of the CW-potential largely get canceled out by the imaginary contributions of eq. (23) and the daisy-resummation term Delaunay:2007wb , which resums the massless modes. Since we can not consistently include the daisy-resummation term, we inevitably have to deal with a complex potential. In the literature, different strategies are chosen to tackle this complication, for example, taking the real part of the potential Delaunay:2007wb or the absolute values of the masses Athron:2022jyi . Unfortunately, these treatments lack a robust theoretical motivation and they lead to unphysical kinks in the potential (see Figure 1 of Profumo:2007wc ), and consequently in the evolution of the background field. Imaginary parts in the potential can be avoided altogether by adhering to a strict perturbative expansion around the leading order minimum of the effective potential Fukuda:1975di ; Nielsen:1975fs ; Gould:2021ccf ; Ekstedt:2022zro ; Gould:2023ovu . This approach requires a careful tracking of the temperature- and field-dependent masses of the fields, and the construction of an appropriate effective field theory for the (light) fields taking part in the PT. This more precise, but also more elaborate approach would be necessary to obtain reliable predictions when we expect an FOPT to occur. As we are currently still investigating whether an FOPT occurs, we will resort to the somewhat ad hoc approach of plugging in the absolute values of the mass into eq. (22). We find this approach to be numerically better behaved than taking the real part, as the evolution of the Higgs field, which is our main interest, does not suffer from large numerical (unphysical) jumps.
@@
Let us first focus on the PT of the Higgs field. By inspecting all 140 BM points, and focusing on temperatures where the couplings are perturbative, we find no hints of an FOPT. A representative evolution of the vevs of the fields as a function of the temperature is illustrated by figure 4. The vevs are found by using the function NMinimize in Mathematica.
@@
Figure 4: Evolution of the field expectation values in the minimum of the potential for a representative third BM point in Table 2. The Higgs field is represented by gray solid, $\phi_{2}$ by dashed pink, and $\phi_{3}$ by dotted cyan.
@@
Our results are in line with, and complement the results of Cline:2009sn , where the EWPT in the singlet-majoron model was studied. In an extensive parameter scan, the authors find that, for the parameter space corresponding to SM with a single additional singlet, there are no strong first-order phase transitions with a zero-temperature non-zero singlet vev. Note that ref.Carena:2019une does find a region with an FOPT and a non-zero singlet vev at zero temperature. Note that the singlet is light in this study, which is therefore not in tension with our results.
@@ 5 Conclusions
Previous studies of the electroweak phase transition in BSM models such as the xSM have shown that the chance of the transition being first order – in the SM it is a crossover – is better for large scalar couplings. A first-order phase transition (FOPT) arises if there is a sufficiently large barrier between separate minima of the potential at some moment in the cosmological history. Such a barrier (which may already exist at zero temperature) can be generated and deformed due to high-temperature effects, giving rise to an FOPT. These effects may come from the temperature-dependence of the leading order thermal masses, or via radiative corrections. We have shown analytically in section 4.2, that the first option does not generate the needed barrier if both added scalars attain a non-zero vev at low temperatures. As for a radiatively induced barrier, no analytical proof is available, but our numerical study shows that no FOPT occurs for any of the BM points. Therefore, the $\mathbb{Z}_{2}$ TRSM parameter space for triple Higgs boson production and that of FOPT appear to be mutually exclusive. To have an FOPT, at least one scalar should have a zero field value at low temperatures. In this scenario, resonant di-Higgs boson production is still possible, enhancing the di-Higgs boson cross-section at the LHC, as one would find within the xSM No:2013wsa ; Chen:2014ask ; Dawson:2015haa ; Li:2019tfd .
Lastly, we point out two ways to achieve both an FOPT and enhanced triple Higgs boson production. The first is to add additional scalar(s) to the $\mathbb{Z}_{2}$ symmetric TRSM. Then, two scalars can have a non-zero vev at low temperatures, contributing to triple Higgs resonance production, and the other(s) can have a zero vev at low temperatures, facilitating an FOPT. Another way is to allow $\mathbb{Z}_{2}$ symmetry-breaking terms to the TRSM. It has been shown that the xSM without $\mathbb{Z}_{2}$-symmetry can exhibit an FOPT while enhancing di-Higgs boson production Kotwal:2016tex ; Alves:2018oct ; Papaefstathiou:2021glr , so similarly, breaking the $\mathbb{Z}_{2}$ symmetry in TRSM would be expected to accommodate the possibility of both an FOPT and triple Higgs boson production enhancement. We leave both of these possibilities for future work.